OpenStudy (anonymous):

simplify (x ^2-10)(x^+10)

OpenStudy (unklerhaukus):

\[(x^2-10)(x^2+10)\\=x^2(x^2+10)-10(x^2+10)\\=\\=\]

OpenStudy (precal):

this is known as a difference of squares

OpenStudy (precal):



OpenStudy (precal):